# Evaluating potential landing sites for the Artemis III mission using a multi-criteria decision making approach

### v1

@@ Abstract
The selection of a landing site within the Artemis Exploration Zone (AEZ) involves multiple factors and presents a complex problem. This study evaluates potential landing sites for the Artemis III mission using a combination of Geographic Information Systems (GIS) and Multi-Criteria Decision Making (MCDM) methodologies, specifically the TOPSIS algorithm. By integrating topographic, illumination, and mineralogy data of the Moon, we assess 1247 locations that meet the Human Landing System (HLS) requirements within 13 candidate regions and Site 004 near the lunar south pole. Criteria considered include surface visibility, HLS-astronaut line of sight, Permanently Shadowed Regions (PSRs), sunlight exposure, direct communication with Earth, geological units, and mafic mineral abundance. Site DM2 (Nobile Rim 2), particularly the point at latitude 84°12’5.61” S and longitude 60°41’59.61” E, is the optimal location for landing. Sensitivity analysis confirms the robustness of our approach, validating the suitability of the best location despite the MCDM method employed and variations in criteria weightings to prioritize illumination and PSRs. This research demonstrates the applicability of GIS-MCDM techniques for lunar exploration and the potential benefits they can bring to the Artemis program.
@@
journal: Journal of  Templates
@@ 1 Introduction
More than 55 years have passed since Neil Armstrong set foot on the Moon in July 1969 during the Apollo 11 mission [1]. Recently, our natural satellite has experienced a resurgence of interest in its exploration, with NASA’s Artemis program as the most clear example of it. Unlike previous missions, which focused on maximizing communication with Earth by landing in the equatorial zone [2, 3, 4, 5, 6], Artemis III aims to explore the lunar south pole region.
@@
The lunar south pole presents a unique opportunity to investigate Permanently Shadowed Regions (PSRs), which may contain reservoirs of water ice. These ice reservoirs are essential for supporting long-term missions on the lunar surface, generating propellants for future missions, and studying trapped volatiles [7]. However, the Artemis Exploration Zone (AEZ), located at 6 latitude from the south pole, features an abrupt and complex topography which may affect astronaut’s Extra-Vehicular Activities (EVA) [8]. Therefore, it is important to identify an area that complies with certain criteria, such as a safe and flat topography to minimize the risk during landing and subsequent operations, direct communication with Earth for efficient data transmission and mission control, as well as adequate exposure to sunlight for solar energy generation and maintaining operational temperatures of equipment, while maximizing the scientific return [9].
@@
Thirteen candidate regions of interest that meet the aforementioned criteria have been officially pre-selected [^1] https://www.nasa.gov/news-release/nasa-identifies-candidate-regions-for-landing-next-americans-on-moon/, see Figure 1. Several studies have been published focusing on the geology and potential astronaut exploration within these sites [10, 11, 12, 13, 14, 15, 16, 17, 18, 19]. However, NASA has not yet selected a landing site for the Artemis III mission [20]. This presents the challenge of selecting the optimal location among several alternatives, where multiple factors or criteria converge. Decision theory provides algorithms and techniques to address such problems, commonly encountered in engineering projects. These are known as Multi-Criteria Decision Making (MCDM) methodologies [21], which we apply in this work to rank the potential landing sites among these regions.
@@
Figure 1: Map of 13 candidate landing regions for Artemis III, in addition to Site 004. Each region measures approximately $\sim$15$\times$15 km. Concentric rings at steps of 1 in latitude.
@@
Since the advent of MCDM approaches in the second half of the 20th century, numerous algorithms and techniques have been developed. Notable among these are the Analytic Hierarchy Process (AHP) [22], Technique for Order of Preference by Similarity to Ideal Solution (TOPSIS) [23], VIseKriterijumska Optimizacija I Kompromisno Resenje (VIKOR) [OpricoviC1998], Elimination and Choice Translating Reality (ELECTRE) [24], Preference Ranking Organization Method for Enrichment Evaluation (PROMETHEE) [25], and DEcision-MAking Trial and Evaluation Laboratory (DEMATEL) [26]. More recent methodologies include Multi-Attributive Border Approximation area Comparison (MABAC) [27], Combined Compromise Solution (CoCoSo) [28], and Additive Ratio Assessment (ARAS) [29]. All these methodologies utilize a decision matrix of alternatives and criteria, akin to the one required for selecting the landing site for the Artemis III mission.
@@
The current availability of high-resolution topographic models of the Moon [30] facilitates the generation of databases of alternatives and criteria for landing site selection. These can be effectively managed using computer-based spatial visualization tools such as Geographic Information Systems (GIS). GIS software is an ideal complement to MCDM methodologies as it allows an easy generation of the decision matrix, which is the starting point of any multi-criteria analysis [31]. In fact, the GIS-MCDM combination has been widely applied to solve various location problems on Earth [32, 33, 34, 35, 36, 37, 38, 39, 40, 41, 42, 43, 44, 45].
@@
The TOPSIS approach, one of the most applied MCDM methodologies, is frequently combined with GIS software for several compelling reasons. Its main advantages include computational simplicity, clear logic, and mathematical robustness [46]. Moreover, it can handle a large number of alternatives based on criteria of diverse nature and different units of measurement. These characteristics make the TOPSIS method particularly attractive when various data layers from GIS software intersect at each pixel identifying the alternatives to be prioritized. The scientific literature demonstrates that the GIS-MCDM combination, previously unused outside the terrestrial environment, is well-suited for solving optimal location problems such as the one posed in this study. Therefore, we employ the GIS-TOPSIS combination to identify the best landing location for the Artemis III mission.
@@ 2 Methods and datasets
The TOPSIS method, developed by [23], is a widely used MCDM methodology in engineering projects [47]. It is based on calculating the relative closeness of each alternative to two fictitious ideal solutions using Euclidean distance: the positive ideal solution and the negative ideal solution. The method asserts that the best alternative is the one closest to the positive ideal solution and farthest from the negative ideal solution. This relative closeness generates a ranking or prioritization of the alternatives. The fundamentals of this MCDM technique are well detailed in the scientific literature [21]. In this work, we apply the TOPSIS method to prioritize the best landing site for the Artemis III mission.
@@
    Step 1. Establish a performance decision matrix: This matrix is composed of rows representing the alternatives $A_{i}~{}(i=1,2,\ldots,m)$ to be evaluated and columns representing the criteria $C_{j}~{}(j=1,2,\ldots,n)$ that influence the assessment process.
@@
    Step 2. Normalize the decision matrix: The decision matrix is normalized using the following expression:
@@
    where $n_{ij}$ is an element of the normalized decision matrix $N$.
@@
    Step 3. Calculate the weighted normalized decision matrix: The normalized decision matrix is multiplied by the global weight vector of the criteria:
@@
    where $w_{j}$ values satisfy $\sum_{j=1}^{n}w_{j}=1$.
@@
    Step 4. Determine the positive ideal $A^{+}$ and negative ideal $A^{-}$ solutions: These solutions are determined by the characteristics of the criteria. For a benefit criterion, the $A^{+}$ solution is the maximum valuation, and for a cost criterion, it is the minimum valuation. The inverse applies for $A^{-}$:
    $A^{+}=\left\{v^{+}_{1},\ldots,v^{+}_{n}\right\}=\left\{\begin{array}[]{ll}max_{i}\left\{v_{ij},j\in J\right\},&i=1,2,\ldots,m\\ min_{i}\left\{v_{ij},j\in J^{\prime}\right\},&i=1,2,\ldots,m\\ \end{array}\right.$ (3)
    $A^{-}=\left\{v^{-}_{1},\ldots,v^{-}_{n}\right\}=\left\{\begin{array}[]{ll}min_{i}\left\{v_{ij},j\in J\right\},&i=1,2,\ldots,m\\ max_{i}\left\{v_{ij},j\in J^{\prime}\right\},&i=1,2,\ldots,m\\ \end{array}\right.$ (4)
    where $J$ corresponds to benefit criteria and $J^{\prime}$ to cost criteria.
@@
    Step 5. Calculate the separation measures of each alternative: The separation distances from $A^{+}$ ( $d^{+}_{i}$ ) and $A^{-}$ ( $d^{-}_{i}$ ) are calculated using the n-dimensional Euclidean distance method:
@@
    Step 6. Calculate the relative closeness to the ideal solution: The closeness coefficient for each alternative is calculated as:
@@
    Step 7. Rank the preference order: Alternatives are ranked in descending order based on $R_{i}$ values. An alternative $A_{i}$ closer to $A^{+}$ and farther from $A^{-}$ has a higher ranking as $R_{i}$ approaches 1.
Regarding the data, we employ 14 Digital Elevation Models (DEMs) of the south pole with a resolution of 5 m/pixel, derived from the Lunar Reconnaissance Orbiter’s (LRO) Lunar Orbiter Laser Altimeter (LOLA) [48]. These account for the 13 candidate regions in addition to Site 004, which we include because of its scientific interest [49]. Solar illumination, PSR, and Earth sunlight reflection maps of 240 m/pixel are also used [50]. We consider as well a 1 $\mu$m absorption depth band map of 1.4 km/pixel from M3 data [51] and a 1:5,000,000 scale geologic map derived from SELENE Kaguya terrain camera stereo (60 m/pix) and LOLA altimetry (100 m/pix) [52]. To handle these maps, we utilize two GIS tools: ArcGIS [53] and Global Mapper [54]. We utilize Bresenham’s line algorithm to generate visibility maps considering the lunar curvature, which include one map representing the total area visible from the HLS windows and another for the line of sight of astronaut-HLS within the 2 km radius limited by the mission requirements [55].
The starting point involves analyzing LOLA topography images [48] of the 14 regions, where potential lunar landing locations will be evaluated and prioritized. This analysis begins by defining the parameters established for the Artemis III HLS [56], which include a landing precision of 100 m from any target landing site and slope requirements with vertical orientations ranging between 0 and 5. We generate a buffer zone around all pixels with a slope of less than 5 in the LOLA topography maps. This ensures that if a lunar module lands on any pixel within the delineated polygons, it will have a safety margin of at least 100 m up to a pixel with a slope greater than 5.
@@ 3.1 Generation of the potential landing locations (alternatives)
Figure 2: Left: All polygons that meet HLS landing requirements for slope compliance. Right: Example of the polygons and pixel point (alternatives; landing locations) within the DM2 site.
@@ 3.2 Criteria to evaluate the potential landing locations
To enhance the geological return of the mission, this criterion evaluates the number of distinct geological units within a 2 km radius of the landing site. This assessment determines the variety of geological units that astronauts can explore during an EVA. The evaluation utilizes the Unified Geological Map of the Moon at a 1:5M scale [52]. Figure 6 illustrates the map of geological units for the Astronaut Exploration Zone (AEZ).
@@
On the Moon, the rich basalts that make up the surface of lunar maria are a primary source of mafic minerals, containing high proportions of magnesium and iron. In the context of lunar exploration and in situ resource utilization, these minerals are particularly valuable due to their composition, which includes elements useful for various applications. For example, mafic materials with higher olivine and pyroxene content melt at lower temperatures, simplifying processing, and they also contain more potentially extractable metallic phases. The strongest mafic signatures are associated with the South Pole-Aitken Basin, indicating the presence of impact melting and basin ejecta that can be sampled and characterized during the Artemis III mission [58]. Utilizing data from the Moon Mineralogy Mapper on band depths of 1 $\mu$m, it is possible to characterize mafic mineral abundances, as this absorption band is sensitive to the presence of pyroxene and olivine basalts [58]. Figure 7 provides the mafic mineral signature map for the AEZ.
Figure 7: Mafic mineral signature in the AEZ. Concentric rings at steps of 1 in latitude.

### v2

@@ Abstract
The selection of a landing site within the Artemis Exploration Zone (AEZ) involves multiple factors and presents a complex problem. This study evaluates potential landing sites for the Artemis III mission using a combination of Geographic Information Systems (GIS) and Multi-Criteria Decision Making (MCDM) methodologies, specifically the TOPSIS algorithm. By integrating topographic, illumination, and mineralogy data of the Moon, we assess 1247 locations that meet the Human Landing System (HLS) requirements within 13 candidate regions and Site 004 near the lunar south pole. Criteria considered include surface visibility, HLS-astronaut line of sight, Permanently Shadowed Regions (PSRs), sunlight exposure, direct communication with Earth, geological units, and mafic mineral abundance. Site DM2 (Nobile Rim 2), particularly the point at latitude 84°12’5.61” S (-84.20156°) and longitude 60°41’59.61” E (60.69989°), is the optimal location for landing. Sensitivity analysis confirms the robustness of our approach, validating the suitability of the best location despite the MCDM method employed and variations in criteria weightings to prioritize illumination and PSRs. This research demonstrates the applicability of GIS-MCDM techniques for lunar exploration and the potential benefits they can bring to the Artemis program.
@@
journal: Acta Astronautica
@@ 1 Introduction
Neil Armstrong first walked on the Moon in July 1969 during the Apollo 11 mission [1]. Over the past 55 years, interest in lunar exploration has grown, particularly with NASA’s Artemis program. Unlike earlier missions that prioritized communication with Earth by landing in the equatorial zone of the lunar nearside [2, 3, 4, 5, 6], Artemis III intends to explore the lunar south pole.
@@
The south pole of the Moon offers a unique opportunity to study Permanently Shadowed Regions (PSRs), which might contain water ice. These ice reserves will enable extended lunar missions, producing propellants for future explorations, and analyzing trapped volatiles [7]. However, the Artemis Exploration Zone (AEZ), situated 6 latitude from the south pole, has a rugged and complex terrain that could impact astronaut Extra-Vehicular Activities (EVA) [8, 9]. Hence, selecting an area that meets specific criteria is vital. These criteria include safe and level topography to reduce risks during landing and operations, direct communication with Earth for efficient data transmission and mission control, adequate sunlight for solar power generation, and maintaining operational temperatures for equipment, all while maximizing scientific returns [10].
@@
Thirteen candidate regions of interest for Artemis III that meet the aforementioned criteria have been officially pre-selected [^1] https://www.nasa.gov/news-release/nasa-identifies-candidate-regions-for-landing-next-americans-on-moon/, see Figure 1. Several studies have been published focusing on the geology and potential astronaut exploration within these sites [11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21]. However, NASA has not yet selected a landing site for the Artemis III mission [22]. This challenge involves selecting the optimal location among several alternatives, each influenced by multiple factors or criteria. Decision theory offers algorithms and techniques to address such problems. These are known as Multi-Criteria Decision Making (MCDM) methodologies [23], which we apply in this work to rank the potential landing sites among these regions.
@@
Figure 1: Map of 13 candidate landing regions for Artemis III, in addition to Site 004. Each region measures approximately $\sim$15$\times$15 km. Terrain farther from the south pole, like Haworth crater, appears larger than it is due to the projection effect. Contours represent 1 km elevation intervals. Concentric rings at steps of 1 in latitude.
@@
Since the introduction of MCDM methods in the latter half of the 20th century, various algorithms and techniques have been developed. Key examples include the DEcision-MAking Trial and Evaluation Laboratory (DEMATEL) [24], Analytic Hierarchy Process (AHP) [25], Technique for Elimination and Choice Translating Reality (ELECTRE) [26], Order of Preference by Similarity to Ideal Solution (TOPSIS) [27], Preference Ranking Organization Method for Enrichment Evaluation (PROMETHEE) [28], VIseKriterijumska Optimizacija I Kompromisno Resenje (VIKOR) [29]. Recently developed approaches include Multi-Attributive Border Approximation area Comparison (MABAC) [30], Combined Compromise Solution (CoCoSo) [31], and Additive Ratio Assessment (ARAS) [32]. Each of these methodologies employs a decision matrix of alternatives and criteria, similar to the matrix needed for selecting the Artemis III mission landing site.
@@
The current availability of detailed topographic models of the Moon [33] facilitates the generation of databases of alternatives and criteria for landing site selection. These can be effectively managed using computer-based spatial visualization tools such as Geographic Information Systems (GIS). GIS software is an ideal complement to MCDM methodologies as it allows an easy generation of the decision matrix, which is the starting point of any multi-criteria analysis [34]. In fact, the GIS-MCDM combination has been widely applied to solve various location problems on Earth [35, 36, 37, 38, 39, 40, 41, 42, 43, 44, 45, 46, 47, 48].
@@
The TOPSIS method is often integrated with Geographic Information Systems (GIS) for several reasons. Its primary benefits include ease of computation, straightforward logic, and mathematical robustness [49]. Additionally, it can manage numerous alternatives based on criteria of varying nature and units of measurement. These attributes make TOPSIS particularly effective when various data layers from GIS software intersect at each pixel, identifying the alternatives to prioritize. The scientific literature indicates that the GIS-MCDM combination, although previously unused beyond terrestrial applications, is well-suited for addressing optimal location problems like the one in this study. Consequently, we utilize the GIS-TOPSIS combination to determine the optimal landing site for the Artemis III mission.
@@ 2 Methods and datasets
The TOPSIS method, introduced by [27], is a popular MCDM technique used in engineering projects [50]. This approach calculates the relative closeness of each alternative to two hypothetical ideal solutions using Euclidean distance: the positive ideal solution and the negative ideal solution. According to TOPSIS, the optimal alternative is the one nearest to the positive ideal solution and furthest from the negative ideal solution. This relative closeness results in a ranking or prioritization of the alternatives. The principles of this MCDM technique are comprehensively detailed in the scientific literature [23]. In this work, we apply the TOPSIS method to prioritize the best landing site for the Artemis III mission.
@@
    Step 1. Construct a performance decision matrix composed of rows representing the alternatives $A_{i}~{}(i=1,2,\ldots,m)$ to be evaluated and columns representing the criteria $C_{j}~{}(j=1,2,\ldots,n)$ that influence the assessment process.
@@
    Step 2. Normalize the decision matrix using the following expression:
@@
    where $n_{ij}$ represents elements in the normalized decision matrix $N$.
@@
    Step 3. Compute the weighted normalized decision matrix by multiplying the normalized decision matrix by the criteria weight vector:
@@
    where $w_{j}$ values are such that $\sum_{j=1}^{n}w_{j}=1$.
@@
    Step 4. Identify the positive ideal $A^{+}$ and negative ideal $A^{-}$ solutions, which depend on the nature of the criteria; for a benefit criterion, $A^{+}$ is the highest value, and for a cost criterion, it is the lowest, and the opposite applies for $A^{-}$:
    $A^{+}=\left\{v^{+}_{1},\ldots,v^{+}_{n}\right\}=\left\{\begin{array}[]{ll}\max_{i}\left\{v_{ij},j\in J\right\}\\ \min_{i}\left\{v_{ij},j\in J^{\prime}\right\}\\ \end{array}\right.$ (3)
    $A^{-}=\left\{v^{-}_{1},\ldots,v^{-}_{n}\right\}=\left\{\begin{array}[]{ll}\min_{i}\left\{v_{ij},j\in J\right\}\\ \max_{i}\left\{v_{ij},j\in J^{\prime}\right\}\\ \end{array}\right.$ (4)
    where $J$ corresponds to benefit criteria and $J^{\prime}$ to cost criteria, and for $i=1,2,\ldots,m$ and $j=1,2,\ldots,n$.
@@
    Step 5. Determine the separation measures of each alternative by calculating the separation distances from $A^{+}$ ( $d^{+}_{i}$ ) and $A^{-}$ ( $d^{-}_{i}$ ) using the n-dimensional Euclidean distance method:
@@
    Step 6. Calculate the relative closeness to the ideal solution by determining the closeness coefficient for each alternative:
@@
    Step 7. Rank the alternatives in descending order based on $R_{i}$ values, where an alternative $A_{i}$ closer to $A^{+}$ and further from $A^{-}$ will have a higher ranking as $R_{i}$ approaches 1.
Regarding the data, the 5 m/pixel Digital Elevation Models (DEMs) derived from LOLA data [51] were chosen for their high spatial resolution [^2] https://pgda.gsfc.nasa.gov/products/78, which is crucial for analyzing terrain features and visibility across the 13 candidate regions and the scientifically significant Site 004 [52]. The slope maps accompanying these DEMs provide essential information for assessing lander stability and traverse planning. The solar illumination, PSRs, and Earth sunlight reflection maps [53], with a resolution of 240 m/pixel, were included to evaluate potential landing sites with regard to energy availability and volatile retention [^3] https://imbrium.mit.edu/BROWSE/EXTRAS/ILLUMINATION/. The 1 $\mu$m absorption depth band map, obtained from M3 data at 1.4 km/pixel [21], was selected to provide insights into mineralogy of the explorable area. Lastly, the 1:5,000,000 scale geologic map, derived from SELENE Kaguya and LOLA data [54], was used to ensure comprehensive geologic context, with spatial resolutions of 60 m/pixel and 100 m/pixel, respectively.
To handle these maps, we utilize two GIS tools: ArcGIS [55] and Global Mapper [56]. We utilize Bresenham’s line algorithm to generate visibility maps considering the lunar curvature, which include one map representing the total area visible from the HLS windows and another for the line of sight of astronaut-HLS within the 2 km radius limited by the mission requirements [57].
The starting point involves analyzing LOLA topography images [51] of the 14 regions, where potential lunar landing locations will be evaluated and prioritized. This analysis begins by defining the parameters established for the Artemis III HLS [58], which include a landing precision of 100 m from any target landing site and slope requirements with vertical orientations ranging between 0 and 5. We generate a buffer zone around all pixels with a slope of less than 5 in the LOLA topography maps. This ensures that if a lunar module lands on any pixel within the delineated polygons, it will have a safety margin of at least 100 m up to a pixel with a slope greater than 5. We note that the reported geolocation uncertainty is approximately 10–20 cm horizontally and 2–4 cm vertically per pixel. However, that around 90% of the 5 m/pixel data is interpolated, meaning it does not consist of directly measured values.
@@ 3.1 Generation of the potential landing locations (alternatives)
Figure 2: Left: All polygons that meet HLS landing requirements for slope compliance. Contours represent 1 km elevation intervals. Concentric rings at steps of 1 in latitude. Right: Example of the polygons and pixel point (alternatives; landing locations) within the DM2 site.
@@ 3.2 Criteria to evaluate the potential landing locations
There are few studies that specifically analyze the selection criteria for lunar exploration missions, and those that exist differ not only in their study area but also in the mission focus (e.g., crewed, robotic rovers). A brief review of the existing literature addressing these criteria follows.
In some cases, a single criterion dominates the decision-making process for selecting high-priority lunar landing sites. For instance, [59] identified the Schrödinger basin as a prime location due to the wide range of science and exploration objectives that could be achieved in this area [60].
For missions targeting the rim of the lunar South Pole-Aitken basin, multiple parameters are considered. These include overall illumination percentage, temporal ground station contact, the size of the area with favorable conditions, and proximity to key basin features [61]. Other studies in the same region emphasize scientific criteria, such as the diversity of rock types representing large areas, the presence of crystalline impact-melt rocks, and basalt flow deposits [62].
Additional parameters, such as the stability of volatiles, PSRs, slope, and temperature conditions, have been used to prioritize landing sites. For example, [63] identified five regions near the lunar north pole and seven near the south pole with high scientific potential based on these criteria.
Common criteria for lunar polar rover missions include a combination of sunlight availability, traversable terrain, subsurface volatiles, and communication capabilities. The selection of Haworth Crater exemplifies this approach, as it offers a favorable intersection of these factors [64]. Similarly, in the Sverdrup-Henson crater region, criteria such as terrain slope, water ice presence, energy sources, and communication capabilities with Earth were considered for site selection [65]. Recently, studies have identified landing areas at the lunar south pole with smooth topography, favorable illumination, moderate temperatures, and volatile material presence as key features for dynamic exploration missions involving rovers or flybys [66]. In addition to NASA, other countries as China has also analyzed the selection of landing sites for their series of Chang’ e missions. Scientific and engineering constraints such as potential for resources utilization, gravity, magnetic field and other environmental features, topographic slope, terrain obstacles, communication ability and temperature, among others, were the main criteria which had influence in the assessment of landing sites [67].
For the Artemis III mission, candidate regions must not only meet HLS requirements but also satisfy additional criteria such as Earth communication capabilities, engineering feasibility, and adequate illumination [68]. These and other relevant criteria are incorporated into our multi-criteria analysis, each of which will be detailed in the following sections.
@@
To enhance the geological return of the mission, this criterion evaluates the number of distinct geological units within a 2 km radius of the landing site. This assessment determines the variety of geological units that astronauts can explore during an EVA. The evaluation utilizes the Unified Geological Map of the Moon at a 1:5M scale [54]. Figure 6 illustrates the map of geological units for the AEZ.
@@
While basalts rich in mafic minerals, such as those containing high proportions of magnesium and iron, are prevalent in lunar maria, the South Pole-Aitken basin is a significant source of mafic materials near the south pole, offering a distinct mafic signature in this region with high proportions of magnesium and iron. In the context of lunar exploration and in situ resource utilization, these minerals are particularly valuable due to their composition, which includes elements useful for various applications. For example, mafic materials with higher olivine and pyroxene content melt at lower temperatures, simplifying processing, and they also contain more potentially extractable metallic phases. The most prominent mafic signatures are found in the South Pole-Aitken Basin, indicating the presence of impact melting and basin ejecta that can be sampled and characterized during the Artemis III mission [21]. Utilizing data from the Moon Mineralogy Mapper on 1 $\mu$m band depths, it is possible to characterize mafic mineral abundances, as this absorption band is sensitive to the presence of pyroxene and olivine in a variety of lithologies, including those likely present in the South Pole-Aitken basin ejecta [21]. Figure 7 provides the mafic mineral signature map for the AEZ.
Figure 7: Mafic mineral signature in the AEZ. Data from [21]. Concentric rings at steps of 1 in latitude.